Kutztown University Department of Mathematics

Colloquium

3:30 p.m. September 23, 2008 Boehm Hall 260

Real Statistics in the Real World

(REALLY?)

Professor James Bobeck

Abstract

What are applied statisticians? What kind of work do they do? Do they sit in a dusty cubicle in the basement and have people give them reams of numbers to add, subtract, multiply, and divide? Nah, that′s work for an accountant or a bookkeeper!

However, if you like thinking, planning, reading publications on diverse topics, talking to people in many areas of technical expertise, (and maybe you do not like arithmetic so much), then perhaps this is a career for you.

This presentation looks at three applied statistics careers and discusses two of them at a high level of detail. It looks into what kind of work you will do, what skill sets you need for success, and how to prepare to get into these careers. Progressing from these high level descriptions, this presentation works down through a few layers of detail. It presents a sampling of examples of projects I have worked on in my careers along with some reasonable statistical models, and some challenges associated with them.

3:00 p.m. refreshments served

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul Ache III, Chair